Alicia recorded the temperature for five days. When he added the five temperatures together he got some of -7.4°C. If he recorded this number is -8.1 -7.4 5.5 6.2 But was missing one record what was in the missing temperature
1
Expert's answer
2020-09-28T09:22:55-0400

Let's denote the missing value as "x". Now we have a linear equation:


"-8.1 - 7.4 + 5.5 + 6.2 + x = -7.4""-3.8 + x = -7.4""x = -7.4 + 3.8 = -3.6"

So, the missing value is "x = -3.6"